OpenStudy (anonymous):

Lt f(x)=x^2+3 and g(x)=x+5. Find g(f(2))-g(3)

OpenStudy (anonymous):

[(2^2+3)+5] - [3+5] = [(4+3)+5] - [8] = (7+5) - 8 = 12 - 8 = 4 I hope this helps!